OpenStudy (anonymous):

Find the discriminant and explain the best method of finding the x-intercepts algebraically for y=-x^2+6x-7

OpenStudy (anonymous):

The quadratic formula is \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] the discriminant is this part \[\sqrt{{b ^{2}-4ac}}\] In this case a= -1 b = 6 and c = -7

OpenStudy (anonymous):

The best method to solve for this problem is to simply the discriminant and solve both the positive and negative solution for x.